In the expression F=k qq/r2 why k=1/4pi epsilon
1
Expert's answer
2019-09-10T13:45:58-0400

Many electrodynamic formulas include a "4\pi" factor. To get rid of it in the practical and the most important formulas, the coefficient k is set equal to "\frac{1}{4\pi \epsilon_0}" The modified formula is thus called rationalized.